Katrina plans to retire in 18 years. She currently has $250,000 in her Individual Retirement Account (IRA), and wants to have $1 million at retirement. What annual interest rate must she earn to reach her goal – assuming she does not save any additional funds?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

1,000,000=250,000*(1+r/100)^18

(1,000,000/250,000)^(1/18)=(1+r/100)

(1+r/100)=1.0801

r=1.0801-1

=8.01%(Approx).